OpenStudy (anonymous):

Please help me in this integration prob. integrate (5 cos^5 (x)) / (sin^7 (x))

myininaya (myininaya):

\[\frac{\cos^5(x)}{\sin^7(x)}=\cot^5(x) \csc^2(x) \\ \text{ recall the following } \\ \frac{d}{dx} \cot(x)=-\csc^2(x) \\ \frac{d}{dx} \csc(x)=-\cot(x) \csc(x) \\ \text{ looks like the first one is going to be the one you want! :)}\] \[\text{ Let } u=\cot(x)\]

OpenStudy (priyar):

Did u get it @pat12345 ?

OpenStudy (anonymous):

so, cot^2 x . csc^2 x . du/csc^2 x = int u^2 du =int (u^3)/3 +c =(cot^3)/3 +C is correct ans ?

myininaya (myininaya):

what happen to the 5th power on the cotangent function ?

OpenStudy (anonymous):

sorry! so, after cancelling scs^2 it will be int -5u^5 du = (-5u^6)/6 +C = -5/6 cot^6 x +C

OpenStudy (anonymous):

Thank you.

myininaya (myininaya):

my notifications aren't working :p but yep that looks amazing! :)